OpenStudy (anonymous):

Sharon purchased a sprinkler system for $1,950 using a six-month deferred payment plan. The interest rate after the introductory period is 23.99%. No down payment is required, but there is a minimum monthly payment of $25. What is the balance at the beginning of the seventh month if only the minimum payment is made each month during the introductory period? a.1950.00 b.2045.91 c.2195.91 (these answers are rounded to the fourth decimal place) I was given this formula, but I can't figure it out: M=P(i(1+i)^nt / (1+i)^nt-1

OpenStudy (anonymous):

The correct answer is b, but I guessed. I would like to know how to get that answer.

OpenStudy (mathstudent55):

Start with 1950 After 1st month: 1950 -25 = 1925 After 2nd month: 1925 - 25 = 1900 After 3rd month: 1900 - 25 = 1875 After 4th month: 1875 - 25 = 1850 After 5th month: 1850 - 25 = 1825 After 6th month: 1825 - 25 = 1800 Future Value = P(1 + i)^n = $1950(1 + 23.99/100/12)^6 = $2195.91 Interest = Future Value - Present value = 2195.91 - 1950 = $245.91 $1800 + $245.91 = $2045.91

Directrix (directrix):

@Iswimstar --> See discussion at http://openstudy.com/study#/updates/510dc843e4b0d9aa3c47828d